OpenStudy (anonymous):

When I put this equation into mathway.com it told me it was an incomplete input. May someone please help? :) Solve This Equation: ____ √x-6 =6 (Lol that's a square root symbol.) a.) -42 b.) 12 c.) 30 d.)42

OpenStudy (usukidoll):

\[\sqrt{x-6}=6\] this ?

OpenStudy (usukidoll):

we can do this manually just square both sides of the equation and solve for x

OpenStudy (usukidoll):

\[(\sqrt{x-6})^2=6^2 \] can you solve it from here?